# Subject-Consistent and Pose-Diverse Text-to-Image Generation

### v1

@@ 1 Introduction
Figure 1: Comparison of subject-consistent generation methods: Vanilla SDXL podell2023sdxl , ConsiStory tewel2024training , StoryDiffusion zhou2024storydiffusion and CoDi (ours). (a&b) Existing methods sacrifice pose diversity for subject consistency, e.g., ConsiStory produces similar poses in Figure 1(a); and the lower right with hands placed in front in Figure 1(b). In contrast, CoDi generates consistent subjects, while matching the pose diversity of Vanilla SDXL. (c) We report two metrics to assess pose quality: 1) fidelity, measuring the distance to the pose of SDXL, and 2) diversity (see Sec.4.1 for details). Our CoDi attains the best performance on both metrics.
@@
While text-to-image (T2I) ramesh2022hierarchical ; saharia2022photorealistic ; rombach2022high ; blattmann2023align models excel in high-quality image generation rombach2022high ; mou2024t2i , they struggle to maintain subject consistency across multiple scenes. Subject-consistent generation (SCG) aims to synthesize images of the same subject across diverse contextual prompts with three key objectives: (1) ensuring subject consistency across generated instances, (2) promoting layout and pose diversity across different instances to avoid repetitive or overly similar compositions, and (3) maintaining prompt fidelity to accurately reflect the semantics of each prompt. The capability enables numerous practical applications including multi-scene narrative for visual storytelling, customizable character design for animation and gaming, and coherent illustration sequences for graphic novels.
@@
Current SCG methods kopiczko2023vera ; ye2023ip primarily rely on training-intensive optimization avrahami2024chosen or mapping networks ruiz2024hyperdreambooth ; gal2023designing to bind subjects to latent representations. These approaches often require computationally expensive fine-tuning per subject or depend on domain-specific encoders, limiting scalability and generalizability. Training-free methods tewel2024training ; zhou2024storydiffusion ; liu2025one have gained significant attention due to their elimination of parameter tuning, strong generalization capabilities, and broad compatibility with diverse diffusion architectures. Current training-free methods—ConsiStory tewel2024training and StoryDiffusion zhou2024storydiffusion —enhance subject consistency by sharing self-attention keys and values across generated images. However, as noted in their limitations tewel2024training ; Hertz_2024_CVPR and evident in Fig. 1, these methods often achieve high consistency at the cost of severely reduced layout and pose diversity, making it challenging to balance all three objectives.
@@
To better balance the three objectives, we propose a training-free framework—subject-Consistent and pose-Diverse generation, dubbed CoDi—that achieves strong subject consistency while preserving diverse poses. Motivated by the progressive nature of diffusion models yue2024exploring —which shows that low-frequency attributes like pose and layout are formed in early denoising steps, while high-frequency details such as facial features emerge later—our CoDi adopts a two-stage strategy: Identity Transport (IT) and Identity Refinement (IR). During the early denoising steps, IT uses optimal transport to align each target image’s features with the reference identity features. Intuitively, this resembles mosaicking: assembling the subject using visual pieces from the reference image, rearranged to match the target pose—thus naturally preserving identity and keeping the original pose. In the later denoising steps, IR further refines subject consistency by guiding each target image to attend to the most salient identity attributes via cross-attention. As shown in Figure 1, our CoDi achieves superior visual results in both subject consistency and pose diversity. We further evaluate pose quality using two metrics: (1) pose fidelity, measuring the distance to the pose of Vanilla SDXL, and (2) pose diversity, quantifying pose variance. CoDi achieves the best performance on both.
@@
We evaluate our method on the existing subject-consistent T2I generation benchmark ConsiStory+ liu2025one . Compared to other training-free approaches, both quantitative and qualitative results validate that our framework achieves better subject consistency while preserving richer layout and pose diversity. It demonstrates a superior trade-off among subject consistency, pose diversity, and prompt fidelity. Further analysis is also provided to demonstrate CoDi’s advantages in pose diversity.
@@ 2 Related Work
To steer text-to-image generation with diffusion models rombach2022high ; podell2023sdxl ; esser2024scaling , various methods have been proposed to incorporate control signals such as depth maps, edge maps, and segmentation mei2025power ; zhang2023adding ; yang2023reco ; Lei_2025_CVPR . Among them, subject consistency (a.k.a identity preservation) has attracted growing attention, aiming to generate a set of images conditioned on a specified subject. Existing subject-consistent generation (SCG) methods can be broadly categorized into two groups: training-based and training-free.
Training-based SCG. Training-based methods require either (1) fine-tuning on additional training data yang2024seed ; li2024photomaker ; li2019storygan ; betker2023improving ; liu2024intelligent or (2) test-time optimization using reference images roich2022pivotal ; gal2023designing ; kumari2023multi ; xiao2024fastcomposer . The first line of work, represented by StoryDALL-E maharana2022storydall and Make-A-Story rahman2023make , incorporates additional modules to capture subject information, followed by fine-tuning on large datasets to enable direct control over the subject given a reference image. The second line of work, exemplified by DreamBooth ruiz2023dreambooth and Textual Inversion gal2022image , optimizes model parameters or token embeddings on the given test images to inject subject identity. Despite their success in maintaining subject consistency, training-based methods suffer from high training costs or significant test-time latency. In contrast, our CoDi is training-free and introduces only mild additional latency.
Training-free SCG. Training-free methods circumvent the need for iterative tuning of model parameters. For instance, 1-Prompt-1-Story liu2025one improves consistency by aligning prompt embeddings across generations. However, textual embedding control alone does not suffice to enforce consistency, often resulting in subject drift. The current leading methods, ConsiStory tewel2024training and StoryDiffusion zhou2024storydiffusion , adopt attention-based mechanisms to promote subject consistency by sharing self-attention keys and values across generated images. However, as noted in their limitation discussions tewel2024training ; Hertz_2024_CVPR , applying attention across a set of images reduces pose diversity. To address this issue, our CoDi explicitly preserves diversity and promotes consistency by aligning early-stage features between the target and reference images via optimal transport.
@@ 3 Method
Our CoDi consists of two stages: Identity Transport (IT) and Identity Refinement (IR). Our IT operates in the early denoising stage to transport identity features from the reference image while preserving the pose and background of the target images. IR is applied in later denoising stages to refine subject consistency in fine-grained details. This two-stage design is inspired by yue2024exploring , which reveals that low-frequency attributes such as pose and layout are determined early in the denoising timesteps, whereas high-frequency components like facial details emerge in later steps. We begin with the setup of subject-consistent generation (SCG), a review of attention-based SCG methods and a brief introduction of optimal transport.
@@ 3.1 Preliminaries
Review of cross-image attention SCG. The current leading training-free SCG methods, ConsiStory tewel2024training and StoryDiffusion zhou2024storydiffusion , adopt attention-based strategies that extend the standard self-attention to cross-image attention mechanism. Formally, let $\{X_{n}\}_{n=1}^{N}$ denote the features of the target images $\{\mathbf{x}_{n}\}_{n=1}^{N}$. For generating $i$-th image, standard self-attention first projects $X_{i}$ to queries $Q_{i}$, keys $K_{i}$, and values $V_{i}$, then compute
@@
While both SCG methods adopt cross-image attention, they differ slightly in implementation: ConsiStory tewel2024training limits attention to masked subject regions, whereas StoryDiffusion zhou2024storydiffusion randomly samples tokens from all regions without subject constraints.
As discussed in their limitations tewel2024training ; Hertz_2024_CVPR , attention-based methods significantly reduce layout diversity. We conjecture that attending to a shared pool of keys and values entangles feature updates across images, implicitly aligning spatial layouts and poses. To mitigate this, prior work tewel2024training introduces components such as attention dropout and query blending. However, these additions increase computational overhead and still fail to recover pose diversity (as shown in Figure 1). In this paper, we draw inspiration from structural learning to simultaneously preserve subject consistency and pose diversity by transporting identity features via optimal transport.
@@
Optimal transport. Optimal transport (OT) villani2008optimal ; monge1781memoire ; zhu2025dynamicmultimodalprototypelearning provides a framework for measuring the distance between two distributions. Specifically, given two sets of support features $\{\mathbf{v}_{m}\}_{m=1}^{M}$ and $\{\mathbf{u}_{n}\}_{n=1}^{N}$, we define two discrete distributions $\mathbb{P}$ and $\mathbb{Q}$ as: [^1] We slightly abuse the notations $\mathbf{x}$ and $N$, which here do not refer to an image or the number of target images.
@@
where $T\in\mathbb{R}^{M\times N}$ is the transport plan, with $T(m,n)\geq 0$ representing the amount of mass moved from $\mathbf{v}_{m}$ to $\mathbf{u}_{n}$, $\langle$,$\rangle$ , ⟩ denotes the Frobenius inner product, $\mathbf{1}_{M}$ is $M$-dimensional all-one vector.
@@ 3.2 Identity Transport
Extract subject features. Masking out background regions offers two benefits for subject consistency: it reduces background interference and computational cost by focusing on the subject alone. We adopt a similar strategy to that of previous methods hertz2022prompt ; tewel2024training , using image-text cross-attention to extract subject masks. Specifically, let $X_{\mathsf{id}}$ denote the features of the reference image $\mathbf{x}_{\mathsf{id}}$ generated from the identity prompt $\mathbf{t}_{\mathsf{id}}$. When generating $\mathbf{x}_{\mathsf{id}}$, we average the cross-attention maps at the final denoising timestep for subject-related tokens (e.g., ‘‘fairy’’), followed by applying Otsu’s method otsu1975threshold to produce a binary mask $M_{\mathsf{id}}$. This mask highlights the subject-relevant regions, from which we extract the subject features as:
@@
With the cost matrix $C$ and the probability masses $\mathbf{a}$ and $\mathbf{b}$, we solve the OT plan $T_{n}$ in Eq. (4) using network simplex algorithm orlin1997polynomial . With the derived $T_{n}$, the subject target features composed by reference subject features are computed as
@@ 3.3 Identity Refinement
Figure 3: Qualitative comparison among Vanilla SDXL podell2023sdxl , ConsiStory tewel2024training , StoryDiffusion zhou2024storydiffusion , and 1-Prompt-1-Story liu2025one . ConsiStory and StoryDiffusion generate similar poses across examples, while 1-Prompt-1-Story preserves pose diversity but struggles with subject consistency. In contrast, our CoDi achieves both.
Table 1: Quantitative comparison of subject consistency, pose diversity and prompt fidelity. Best results are marked in bold.
Method Subject Consistency Pose Prompt
Vanilla SDXL podell2023sdxl 0.8417 0.8010 0.3139 0.0772 0.9082
1-Prompt-1-Story liu2025one 0.8627 0.8233 0.2959 0.0662 0.8814
ConsiStory tewel2024training 0.8751 0.8428 0.2336 0.0621 0.9148
StoryDiffusion zhou2024storydiffusion 0.8776 0.8471 0.2356 0.0605 0.9038
CoDi (ours) 0.8809 0.8514 0.2136 0.0758 0.9041
@@ 4.1 Setup
Benchmark. We evaluate our CoDi on the standard subject-consistent benchmark, ConsiStory+liu2025one , which comprises nearly 200 prompt sets and supports the generation of over 1,100 images. Each prompt set includes a subject described in a specific style, with multiple frame-specific descriptions.
Baselines and implementation details. We compare our CoDi with SoTA training-free SCG methods, including ConsiStory tewel2024training , StoryDiffusion zhou2024storydiffusion and 1-Prompt-1-Story liu2025one . We reproduce all baselines using their officially released code. All methods are implemented using the same backbone model, Stable Diffusion XL 1.0 podell2023sdxl , with an image resolution of $1024\times 1024$, except for StoryDiffusion, which is evaluated at $768\times 768$ due to its high memory consumption, following its original setting. To ensure fairness, identical noise seeds are used for all methods. We set the hyperparameter $\alpha$ in Eq. (10) to select the top 50% of reference features.
Evaluation metrics. Our evaluation framework assesses the quality of generated images from three aspects: (1) subject consistency, (2) pose diversity, and (3) prompt fidelity. Subject consistency is evaluated by computing the average pairwise cosine similarity (or distance) between image embeddings within each target image set. We use three image encoders for this evaluation: CLIP-I hessel2021clipscore , DINO-v2 oquab2023dinov2 , and DreamSim fu2023dreamsim . To evaluate pose diversity, we extract 2D human joint coordinates using ViTPose’s pose estimation model xu2022vitpose . To eliminate global variations in translation, rotation, and scale, we align poses using Procrustes analysis schonemann1966generalized , inspired by standard practices in face alignment 9442331 . The pose diversity score is then computed as the average Euclidean distance between corresponding keypoints across aligned image pairs. A higher score indicates greater pose diversity. For prompt fidelity, we use CLIP-Score hessel2021clipscore to measure the cosine similarity between each image embedding and its corresponding textual prompt embedding. See Appendix B for more details.

### v2

# CoDi: Subject-Consistent and Pose-Diverse Text-to-Image Generation

@@ 1 Introduction
While text-to-image (T2I) Ramesh et al. (2022); Saharia et al. (2022); Rombach et al. (2022); Blattmann et al. (2023) models excel in high-quality image generation Rombach et al. (2022); Mou et al. (2024), they struggle to maintain subject consistency across multiple scenes. Subject-consistent generation (SCG) aims to synthesize images of the same subject across diverse contextual prompts with three key objectives: (1) ensuring subject consistency across generated instances, (2) promoting layout and pose diversity across different instances to avoid repetitive or overly similar compositions, and (3) maintaining prompt fidelity to accurately reflect the semantics of each prompt. The capability enables numerous practical applications including multi-scene narrative for visual storytelling, customizable character design for animation and gaming, and coherent illustration sequences for graphic novels.
@@
Current SCG methods Kopiczko et al. (2024); Ye et al. (2023) primarily rely on training-intensive optimization Avrahami et al. (2024) or mapping networks Ruiz et al. (2024); Gal et al. (2023b) to bind subjects to latent representations. These approaches often require computationally expensive fine-tuning per subject or depend on domain-specific encoders, limiting scalability and generalizability. Training-free methods Tewel et al. (2024); Zhou et al. (2024) have gained significant attention due to their elimination of parameter tuning, strong generalization capabilities, and broad compatibility with diverse diffusion architectures. Current training-free methods—ConsiStory Tewel et al. (2024) and StoryDiffusion Zhou et al. (2024)—enhance subject consistency by sharing self-attention keys and values across generated images. However, as noted in their limitations Tewel et al. (2024); Hertz et al. (2024) and evident in Fig. 1, these methods often achieve high consistency at the cost of severely reduced layout and pose diversity, making it challenging to balance all three objectives.
@@
To better balance the three objectives, we propose a training-free framework—subject-Consistent and pose-Diverse generation, dubbed CoDi—that achieves strong subject consistency while preserving diverse poses. Motivated by the progressive nature of diffusion models Yue et al. (2024)—which shows that low-frequency attributes like pose and layout are formed in early denoising steps, while high-frequency details such as facial features emerge later—our CoDi adopts a two-stage strategy: Identity Transport (IT) and Identity Refinement (IR). During the early denoising steps, IT uses optimal transport to align each target image’s features with the reference identity features. Intuitively, this resembles mosaicking: assembling the subject using visual pieces from the reference image, rearranged to match the target pose—thus naturally preserving identity and keeping the original pose. In the later denoising steps, IR further refines subject consistency by guiding each target image to attend to the most salient identity attributes via cross-attention. As shown in Figure 1, CoDi achieves superior visual results in both subject consistency and pose diversity, and quantitatively demonstrates advantages over existing methods in balancing this trade-off.
@@
We evaluate our method on the existing T2I SCG benchmark ConsiStory+ Liu et al. (2025). Compared to other training-free approaches, both quantitative and qualitative results validate that our framework achieves better subject consistency while preserving richer layout and pose diversity. It demonstrates a superior trade-off among subject consistency, pose diversity, and prompt fidelity. Further analysis is also provided to demonstrate CoDi’s advantages in pose diversity.
Figure 1: Comparison of subject-consistent generation methods: Vanilla SDXL Podell et al. (2023), ConsiStory Tewel et al. (2024), StoryDiffusion Zhou et al. (2024) and CoDi (ours). (a&b) Existing methods sacrifice pose diversity for subject consistency, e.g., ConsiStory produces similar poses in Figure 1 (a); and the lower right with hands placed in front in Figure 1 (b). In contrast, CoDi generates consistent subjects, while matching the pose diversity of Vanilla SDXL. (c) Subject consistency vs. pose diversity. Current methods struggle to balance the two, whereas CoDi achieves both effectively.
@@ 2 Related Work
To steer T2I generation with diffusion models Rombach et al. (2022); Podell et al. (2023); Nan et al. (2025); Esser et al. (2024); Zhu et al. (2025a); Wang et al. (2025a); Hu et al. (2023); Lin et al. (2025a); Zhou et al. (2025); Lin et al. (2025b); Wang et al. (2025b), various methods have been proposed to incorporate control signals such as depth maps, edge maps, and segmentation Mei et al. (2025); Zhang et al. (2023); Yang et al. (2023); Lei et al. (2025); Chen et al. (2025). Among them, subject consistency (a.k.a identity preservation) has attracted growing attention, aiming to generate a set of images conditioned on a specified subject. Existing subject-consistent generation (SCG) methods can be broadly categorized into two groups: training-based and training-free.
Training-based SCG. Training-based methods require either (1) fine-tuning on additional training data Yang et al. (2024); Li et al. (2024; 2019); Betker et al. (2023); Liu et al. (2024) or (2) test-time optimization using reference images Roich et al. (2022); Gal et al. (2023b); Kumari et al. (2023); Xiao et al. (2024). The first line of work, represented by StoryDALL-E Maharana et al. (2022) and Make-A-Story Rahman et al. (2023), incorporates additional modules to capture subject information, followed by fine-tuning on large datasets to enable direct control over the subject given a reference image. The second line of work, exemplified by DreamBooth Ruiz et al. (2023) and Textual Inversion Gal et al. (2023a), optimizes model parameters or token embeddings on the given test images to inject subject identity. Despite their success in maintaining subject consistency, training-based methods suffer from high training costs or significant test-time latency. In contrast, our CoDi is training-free and introduces only mild additional latency.
Training-free SCG. Training-free methods circumvent the need for iterative tuning of model parameters. For instance, 1Prompt1Story Liu et al. (2025) improves consistency by aligning prompt embeddings across generations. However, textual embedding control alone is insufficient to to enforce consistency, often resulting in subject drift. The current leading methods, ConsiStory Tewel et al. (2024) and StoryDiffusion Zhou et al. (2024), adopt attention-based mechanisms to promote subject consistency by sharing self-attention keys and values. However, as noted in their limitation discussions Tewel et al. (2024); Hertz et al. (2024), applying attention across a set of images reduces pose diversity. To address this issue, our CoDi explicitly preserves diversity and promotes consistency by aligning early-stage features between the target and reference images via optimal transport.
@@ 3 Method
Our CoDi consists of two stages: Identity Transport (IT) and Identity Refinement (IR). Our IT operates in the early denoising stage to transport identity features from the reference image while preserving the pose and background of the target images. IR is applied in later denoising stages to refine subject consistency in fine-grained details. This two-stage design is inspired by Yue et al. (2024), which reveals that low-frequency attributes such as pose and layout are determined early in the denoising timesteps, whereas high-frequency components like facial details emerge in later steps. We begin with the setup of subject-consistent generation (SCG), a review of attention-based SCG methods and a brief introduction of optimal transport.
@@ 3.1 Preliminaries
Review of cross-image attention SCG. The current leading training-free SCG methods, ConsiStory Tewel et al. (2024) and StoryDiffusion Zhou et al. (2024), adopt attention-based strategies that extend the standard self-attention to cross-image attention mechanism. Formally, let $\{X_{n}\}_{n=1}^{N}$ denote the features of the target images $\{\mathbf{x}_{n}\}_{n=1}^{N}$. For generating $i$-th image, standard self-attention first projects $X_{i}$ to queries $Q_{i}$, keys $K_{i}$, and values $V_{i}$, then compute
@@
While both SCG methods adopt cross-image attention, they differ slightly in implementation: ConsiStory Tewel et al. (2024) limits attention to masked subject regions, whereas StoryDiffusion Zhou et al. (2024) randomly samples tokens from all regions without subject constraints.
As discussed in their limitations Tewel et al. (2024); Hertz et al. (2024), attention-based methods significantly reduce layout and pose diversity. We conjecture that attending to a shared pool of keys and values entangles feature updates across images, implicitly aligning spatial layouts and poses. To mitigate this, prior work Tewel et al. (2024) introduces components such as attention dropout and query blending. However, these additions increase computational overhead and still fail to recover pose diversity (as shown in Figure 1). In this paper, we draw inspiration from structural learning to simultaneously preserve subject consistency and pose diversity by transporting identity features to each target image via optimal transport.
@@
Optimal transport. Optimal transport (OT) Villani et al. (2008); Monge (1781); Zhu et al. (2025b) provides a framework for measuring the distance between two distributions. Specifically, given two sets of support features $\{\mathbf{v}_{m}\}_{m=1}^{M}$ and $\{\mathbf{u}_{n}\}_{n=1}^{N}$, we define two discrete distributions $\mathbb{P}$ and $\mathbb{Q}$ as: [^1] We slightly abuse the notations $\mathbf{x}$ and $N$, which here do not refer to an image or the number of target images.
@@
where $T\in\mathbb{R}^{M\times N}$ is the transport plan, with $T(m,n)\geq 0$ representing the amount of mass moved from $\mathbf{v}_{m}$ to $\mathbf{u}_{n}$, $\langle,\rangle$ denotes the Frobenius inner product, $\mathbf{1}_{M}$ is $M$-dimensional all-one vector.
@@ 3.2 Identity Transport
Extract subject features. Masking out background regions offers two benefits for subject consistency: it reduces background interference and computational cost by focusing on the subject alone. We adopt a similar strategy to that of previous methods Hertz et al. (2023); Tewel et al. (2024), using image-text cross-attention to extract subject masks. Specifically, let $X_{\mathsf{id}}$ denote the features of the reference image $\mathbf{x}_{\mathsf{id}}$ generated from the identity prompt $\mathbf{t}_{\mathsf{id}}$. When generating $\mathbf{x}_{\mathsf{id}}$, we average the cross-attention maps at the final denoising timestep for subject-related tokens (e.g., ‘‘fairy’’), followed by applying Otsu’s method Otsu et al. (1975) to produce a binary mask $M_{\mathsf{id}}$. This mask highlights the subject-relevant regions, from which we extract the subject features as:
@@
With the cost matrix $C$ and the probability masses $\mathbf{a}$ and $\mathbf{b}$, we solve the OT plan $T_{n}$ in Eq. (4) using network simplex algorithm Orlin (1997). With the derived $T_{n}$, the subject target features composed by reference subject features are computed as
@@ 3.3 Identity Refinement
Figure 3: Qualitative comparison among Vanilla SDXL Podell et al. (2023), ConsiStory Tewel et al. (2024), StoryDiffusion Zhou et al. (2024), and 1Prompt1Story Liu et al. (2025). ConsiStory and StoryDiffusion generate similar poses across examples, while 1Prompt1Story preserves pose diversity but struggles with subject consistency. In contrast, our CoDi achieves both.
Table 1: Quantitative comparison of subject consistency, pose diversity and prompt fidelity. Best results are marked in bold. $\uparrow$ indicates higher is better, and $\downarrow$ indicates lower is better.
Method Subject Consistency Pose Prompt CLIP-I ($\uparrow$) DINO-v2 ($\uparrow$) DreamSim ($\downarrow$) Diversity ($\uparrow$) Fidelity ($\uparrow$) Vanilla SDXL Podell et al. (2023) 0.8417 0.8010 0.3139 0.0772 0.9082 1Prompt1Story Liu et al. (2025) 0.8627 0.8233 0.2959 0.0662 0.8814 ConsiStory Tewel et al. (2024) 0.8751 0.8428 0.2336 0.0621 0.9148 StoryDiffusion Zhou et al. (2024) 0.8776 0.8471 0.2356 0.0605 0.9038 CoDi (ours) 0.8809 0.8514 0.2136 0.0758 0.9041
@@ 4.1 Setup
Benchmark. We evaluate our CoDi on the standard SCG benchmark, ConsiStory+Liu et al. (2025), which comprises nearly 200 prompt sets and supports the generation of over 1,100 images. Each prompt set includes a subject described in a specific style, with multiple frame-specific descriptions.
Baselines and implementation details. We compare our CoDi with SoTA training-free SCG methods, including ConsiStory Tewel et al. (2024), StoryDiffusion Zhou et al. (2024) and 1Prompt1Story Liu et al. (2025). We reproduce all baselines using their official released code. All methods are implemented using the same backbone model, Stable Diffusion XL 1.0 Podell et al. (2023), with an image resolution of $1024\times 1024$, except for StoryDiffusion, which is evaluated at $768\times 768$ due to its high memory consumption, following its original setting. To ensure fairness, identical noise seeds are enforced for all methods, ensuring that each prompt is initialized with the same random noise input. Hyperparameter $\alpha$ in Eq. (10) selects the top 50% of reference features.
Evaluation metrics. Our evaluation framework assesses the quality of generated images from three aspects: (1) subject consistency, (2) pose diversity, and (3) prompt fidelity. Subject consistency is evaluated by computing the average pairwise cosine similarity (or distance) between image embeddings within each target image set. We use three image encoders for this evaluation: CLIP-I Hessel et al. (2021), DINO-v2 Oquab et al. (2023), and DreamSim Fu et al. (2023). To evaluate pose diversity, we extract 2D human joint coordinates using ViTPose’s pose estimation model Xu et al. (2022). To eliminate global variations in translation, rotation, and scale, we align poses using Procrustes analysis Schönemann (1966), inspired by standard practices in face alignment Lin et al. (2021). The pose diversity score is then computed as the average Euclidean distance between corresponding keypoints across aligned image pairs. A higher score indicates greater pose diversity. For prompt fidelity, we use CLIP-Score Hessel et al. (2021) to measure the cosine similarity between image and textual prompt embeddings. See Appendix B for more details.